Can you help me with these problems. I have been working on them all day.
Complete the order pairs sto that each is a solution to the given Linear equation.
y=2x+7
(a)(0,?)(b)(4,?)
3x-7y=30
(a)(3,?)(b)(?,3)
Thanks for taking time to look at and answer my question!!!

Answer by Earlsdon(6294)

This is a simple matter of substituting the given values of x or y and solving the equation for the other variable.
1) y = 2x + 7
(a)Given (0, ?), substitute x = 0 into the equation and solve for y.
y = 2(0) + 7
y = 7
The ordered pair for (a) is : (0, 7)
(b) Given (4, ?), substitute x = 4 into the equation and solve for y.
y = 2(4) + 7
y = 8 + 7
y = 15
The ordered pair for (b) is: (4, 15) Get the idea?
I'll do (a) of the second problem, then you try (b).
2) 3x - 7y = 30
(a) Given (3, ?), substitute x = 3 into the equation and solve for y.
3(3) - 7y = 30
9 - 7y = 30 Subtract 9 from both sides.
-7y = 21 Divide both sides by -7
y = -3
The ordered pair for (a) is: (3, -3)
Try (b), but in this one, you are given the y-coordinate (3) and you are to solve for x. Otherwise, the procedure is the same.